OpenStudy (anonymous):

Solve: 2(x + 1) + 4 = 8 Answer 6.5 4 3 1

jhonyy9 (jhonyy9):

2x+2+4=8 2x=8-6 2x=2 x=2/2 x=1

OpenStudy (anonymous):

2x+2+4=8 =>2x+6 =8 =>2x=2 =>x=1